# Simple Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

### v1

@@ Abstract
Vision-language models (VLMs) have achieved remarkable success across diverse tasks by leveraging rich textual information with minimal labeled data. However, deploying such large models remains challenging, particularly in resource-constrained environments. Knowledge distillation (KD) offers a well-established solution to this problem; however, recent KD approaches from VLMs often involve multi-stage training or additional tuning, increasing computational overhead and optimization complexity. In this paper, we propose Dual-Head Optimization (DHO)—a simple yet effective KD framework that transfers knowledge from VLMs to compact, task-specific models in semi-supervised settings. Specifically, we introduce dual prediction heads that independently learn from labeled data and teacher predictions, and propose to linearly combine their outputs during inference. We observe that DHO mitigates gradient conflicts between supervised and distillation signals, enabling more effective feature learning than single-head KD baselines. As a result, extensive experiments show that DHO consistently outperforms baselines across multiple domains and fine-grained datasets. Notably, on ImageNet, it achieves state-of-the-art performance, improving accuracy by 3% and 0.1% with 1% and 10% labeled data, respectively, while using fewer parameters.
@@ 1 Introduction
Figure 1: (Top): Previous multi-stage VLMs distillation methods. (Bottom): Our single-stage distillation framework with DHO.
Vision-Language Models (VLMs), which learn joint vision-language representations through contrastive learning [72, 42], have shown promising results in few-shot scenarios by leveraging rich textual information from pre-training. Recent work has explored various adaptation strategies for VLMs with limited labeled data, including parameter-efficient approaches such as linear probing [72, 51, 40], lightweight adapters [28, 115, 109, 78], and prompt-based fine-tuning methods [43, 120, 119, 44, 121, 45, 117, 75, 113, 49], demonstrating the potential of VLMs for data-limited visual recognition tasks.
However, the substantial computational requirements and large model sizes of foundation models often make them impractical for deployment in resource-constrained environments, such as mobile devices [77]. While compact models could address these challenges, they typically struggle with limited performance, especially with limited labeled data. Semi-supervised learning techniques [80, 2, 7, 118] leveraging unlabeled data can help address these performance limitations, but they are often suboptimal for small models compared to the performance of foundation models, such as VLMs.
@@
Knowledge distillation (KD) [38] has emerged as a promising approach for transferring capabilities from large VLMs to more compact models. However, existing methods primarily focus on distilling large VLMs into smaller ones through general-purpose training [24, 85, 98, 101, 90, 88, 103] or preceding with unsupervised distillation stage [91, 97], which often requires additional fine-tuning for specific target tasks (Figure 1-(Top)). This multi-stage pipeline not only incurs extra computational overhead but also limits the direct transfer of the teacher model’s zero- and few-shot capabilities to the student model for task-specific applications, suggesting a need for more efficient distillation methods that can preserve the teacher’s expertise for the intended task.
In contrast, conventional KD methods—such as logits distillation [38, 13] and feature matching [101]—enable efficient single-stage distillation. However, we find these approaches suboptimal in semi-supervised settings. We attribute this limitation to the discrepancy between the labeled training data and the teacher model’s pre-trained knowledge [116], often leading to gradient conflicts in the feature extractor ( in Figure 4). Such conflicts are well known to hinder effective feature learning [110, 54, 11]. This issue is particularly challenging in few-shot settings, where the strong distillation signal of teacher can overwhelm the limited labeled data, necessitating careful balancing between the two signals.
@@
To address the above issue, we propose a simple yet effective distillation framework, DHO (Dual-Head Optimization), which jointly leverages labeled samples and the teacher model’s probabilistic outputs by learning two distinct heads, each optimized with a separate loss: the supervised loss and the KD loss, respectively (Figure 1–(Bottom)). We observe that DHO mitigates gradient conflicts arising from the two distinct training signals ( in Figure 4). As a result, DHO yields improved feature representations compared to conventional KD baselines, as demonstrated in Section 4.2 and Figure 7. Furthermore, to control the relative influence of the teacher and supervised predictions, we propose to generate the final output by linearly combining the predictions from both heads, whose effectiveness is also empirically validated in Figure 8.
To validate the empirical effectiveness of the proposed DHO, we conduct extensive experiments across 11 different datasets including various tasks, such as generic object recognition [76, 26], fine-grained classification  [67, 59, 69], domain-specific recognition [6], scene understanding [99], texture analysis [17], satellite imagery [34], and human actions [81]. The experimental results validate the effectiveness of DHO over conventional KD methods across all the datasets (Figure 2-(Left)). Furthermore, in comparison to previous state-of-the-art (SoTA) semi-supervised methods on the ImageNet, DHO achieves new SoTA performance, improving accuracy by 3% and 0.1% with 1% and 10% labeled data, respectively, while using fewer parameters (Figure 2-(Right)).
@@
  • We identify gradient conflicts in conventional single-head KD baselines under semi-supervised settings and propose a simple yet effective framework, DHO, which mitigates such conflicts using dual prediction heads that separately learn from labeled data and teacher signals.
  • We show that DHO improves feature learning compared to baselines and supports post-training adjustment of supervised and teacher signals via a linear combination of the dual head outputs.
  • We demonstrate that DHO outperforms single-head KD baselines across 11 datasets, achieving state-of-the-art results on ImageNet in semi-supervised settings with fewer parameters.
2 Related Work
Vision-language pre-training.
The emergence of vision-language pre-training has marked a significant breakthrough, enabling the use of extensive image-text pairs collected from the web [95, 9] to train powerful vision encoders transferable to various vision tasks [27, 114]. Early works such as CLIP [72] and ALIGN [42] leveraged contrastive learning techniques to align images and text into a joint representation space, facilitating zero-shot transfer via language prompts. Building on these foundations, subsequent research has focused on improving vision-language models through enhanced training methodologies [19, 29, 107, 112], as well as scaling models and datasets [107, 53, 18, 83, 15, 22, 84, 30] with their zero-shot transfer capabilities [42, 111, 71, 55]. In contrast, our work focuses specifically on target tasks with compact models, aiming to distill knowledge from these large VLMs effectively.
Data-limited adaptation of VLMs.
To preserve pretrained semantic features of VLMs during adaptation with limited data, several approaches have been proposed. Prompt tuning [50], initially designed for language models, has been successfully extended to vision tasks. Various methods [43, 120, 119, 44, 121, 45, 60, 117, 75, 113, 49] have demonstrated the effectiveness of training learnable prompts while keeping the base model frozen. Adapters [28, 115, 109, 78] provide an alternative approach by introducing lightweight, trainable modules while maintaining the pre-trained backbone intact. LP++ [40] has shown that simple linear layers can effectively adapt CLIP representations in data-limited settings. Note that our work is orthogonal to these approaches: we aim to distill the knowledge of pretrained VLMs into compact models under data-scarce scenarios, making these adaptation methods complementary and applicable to both teacher VLMs in our framework and student models when they are also VLMs.
Knowledge Distillation (KD) [38] enables transferring knowledge from large teacher models to compact student architectures, particularly in data-constrained settings. Researchers have explored synthetic data generation [57, 47, 65, 106, 10, 105, 23, 66, 70, 108, 56, 86, 96], semi-supervised [13, 33, 21, 102], and unsupervised KD using self-supervised teachers [25, 1, 64, 94, 100, 79]. In the VLM domain, recent works [24, 98, 85, 101, 90, 88, 103] distill from large-scale vision-language models into smaller architectures, often using transductive [46, 14] or multi-stage unsupervised strategies [91, 97, 63]. Meanwhile, KD remains challenging due to numerous issues, including model capacity gaps [16, 62, 122, 39, 52] and inconsistencies between soft and hard targets [116]. These challenges are further complicated by misalignment between labeled data and foundational knowledge, especially in few-shot learning scenarios where limited labeled examples may not fully capture the rich semantic understanding of foundation models. While previous KD methods adopt dual-head architectures—such as SSKD [33], which trains separate heads for labeled and unlabeled data, and DHKD [104], which applies a binary KD loss to avoid neural collapse [68]—they do not target distillation from foundation models or combine predictions at inference. In contrast, DHO addresses gradient conflicts in this setting and leverages dual-head aggregation at inference to enhance performance with minimal hyperparameter tuning costs.
3 Methodology
We now elaborate on our approach, Dual-Head Optimization (DHO), a simple yet effective Knowledge Distillation (KD) framework for semi-supervised settings. We first present the preliminaries of our method, such as background on VLMs, problem formulation, single-head KD baselines, in §3.1, followed by our proposed DHO method in §3.2.
3.1 Preliminary
@@
Our work builds upon Vision-Language Models (VLMs) such as CLIP [72] and ALIGN [42]. These models consist of multimodal encoders: an image encoder $f_{\mathbb{X}}:\mathbb{X}\to\mathbb{R}^{d}$ and a text encoder $f_{\mathbb{T}}:\mathbb{T}\to\mathbb{R}^{d}$ where $\mathbb{X}$ and $\mathbb{T}$ denote the domains of images and texts, respectively. The encoders project their respective inputs, i.e., either an image $x\in\mathbb{X}$ or a text description $t\in\mathbb{T}$, into a shared embedding space $\mathbb{R}^{d}$. They are trained via contrastive learning [12] to align corresponding image-text pairs in the shared embedding space.
For zero-shot classification of VLMs across $C$ classes, we use predefined prompt templates, e.g., “a photo of a [CLASS]”, where [CLASS] is the name of class. Given a set of $C$ target classes, i.e., $y\in[C]:=\{1,\ldots,C\}$, we generate prompted text descriptions $\{t_{1},t_{2},\dots,t_{C}\}$. The zero-shot inference for categorical probability vector $p$ over the $C$ classes is defined as follows:
$p=\sigma\left(\frac{1}{\zeta}[\mathtt{CosSim}(f_{\mathbb{X}}(x),f_{\mathbb{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathbb{X}}(x),f_{\mathbb{T}}(t_{C}))]^{\top}\right)\in\Delta^{C-1},$ (1)
where $\Delta^{C-1}$ is the probability simplex of dimension $C-1$, $\sigma:\mathbb{R}^{C}\to\Delta^{C-1}$ denotes the softmax function defined as $\sigma(l)=\left[\frac{\mathtt{exp}(l_{1})}{\sum_{c=1}^{C}\mathtt{exp}(l_{c})},\ldots,\frac{\mathtt{exp}(l_{C})}{\sum_{c=1}^{C}\mathtt{exp}(l_{c})}\right]^{\top}$, $\zeta\in\mathbb{R}_{>0}$ is the temperature scaling [38], and $\mathtt{CosSim}(x,y)=\frac{x^{\top}y}{||x||_{2}\cdot||y||_{2}}$ is the cosine similarity function. The final classification is determined by $\arg\max_{c\in[C]}p_{c}$. Furthermore, for few-shot transfer of VLMs, we adopt Tip-Adapter [115], which builds a key-pair queue using few labeled examples.
@@
In this work, we focus on transferring knowledge from VLMs to compact, task-specific models under few-shot or low-shot semi-supervised learning scenarios, where both labeled and unlabeled data are utilized. Specifically, given a $K$-shot and $C$-class classification problem, we are provided with a labeled dataset $\mathcal{D}^{(l)}=\{(x^{(l)}_{n},y_{n})\}_{n=1}^{N}$, where $N=K\times C$ is the total number of labeled examples, and $y_{n}\in[C]$ denotes the class labels. Additionally, we have access to an unlabeled dataset $\mathcal{D}^{(u)}=\{x^{(u)}_{m}\}_{m=1}^{M}$ consisting of $M$ unlabeled images. Low-shot learning represents a more realistic setting than traditional few-shot learning where only a small fraction, e.g., 1% ($\frac{N}{N+M}\approx 0.01$) or 10% ($\frac{N}{N+M}\approx 0.1$) of the total dataset is labeled. Our goal is then to develop an student model by leveraging $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$, guided by the knowledge of the VLM encoders $f_{\mathbb{X}}$ and $f_{\mathbb{T}}$. The student model consists of a feature extractor $g:\mathbb{X}\to\mathbb{R}^{d}$ and a prediction head $h:\mathbb{R}^{d}\to\mathbb{R}^{C}$, followed by the softmax function $\mathtt{\sigma}:\mathbb{R}^{C}\to\Delta^{C-1}$.
Figure 3: Conceptual illustration on KD frameworks, Single-Head Optimization (SHO) and Dual-Head Optimization (DHO), for semi-supervised settings. As demonstrated in Figure 4, we observe gradient conflict of SHO. In contrast, DHO mitigates such conflicts by leveraging dual-head architectures.
Single-head KD baselines.
Building upon traditional KD approaches [38], numerous studies have explored semi-supervised KD [13, 33, 21, 102], unsupervised KD using self-supervised teacher models [25, 1, 64, 94, 100, 79], and KD from VLMs [24, 98, 85, 101, 90, 88, 103], to name just a few. Our method builds on simple logit distillation in semi-supervised settings [38, 13] that combines supervised loss $\mathcal{L}_{\text{CE}}$ on the labeled dataset $\mathcal{D}^{(l)}$ with KD loss $\mathcal{L}_{\text{KD}}$ on both labeled and unlabeled datasets $\mathcal{D}^{(l)}\cup\mathcal{D}^{(u)}$, i.e., $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, where $\lambda\in[0,1]$ is a loss balancing hyperparameter. Specifically, the supervised loss $\mathcal{L}_{\text{CE}}$ and KD loss $\mathcal{L}_{\text{KD}}$ are defined as follows:
@@
$\mathcal{L}_{\text{KD}}$ $=\frac{1}{N}\sum_{n}D_{\mathrm{KL}}\left[p_{n}^{(l)}\|\sigma(h(z_{n}^{(l)}))\right]+\frac{1}{M}\sum_{m}D_{\mathrm{KL}}\left[p_{m}^{(u)}\|\sigma(h(z_{m}^{(u)}))\right],$ (3)
where $\ell$ denotes $D_{\mathrm{KL}}$ represent cross-entropy and Kullback-Leibler divergence, respectively. $z_{n}^{(l)}=g(x_{n}^{(l)})$ and $z_{m}^{(u)}=g(x_{m}^{(u)})\in\mathbb{R}^{d}$ are feature representations obtained by the feature extractor $g$. $p_{n}^{(l)}$ and $p_{m}^{(u)}\in\Delta^{C-1}$ are the categorical probability vectors of labeled $x_{n}^{(l)}$ and unlabeled data $x_{m}^{(u)}$, respectively, obtained by teacher VLM encoders $f_{\mathbb{X}}$ and $f_{\mathbb{T}}$, as described in Eq. 1. To optimize the parameters of $g$ and $h$, we use a mini-batch version of the above objective with stochastic gradient descent. Another well-studied single-head KD baseline is feature distillation, which leverages mean squared error (MSE) loss to directly align feature representations of student $g$ and teacher $f_{\mathbb{X}}$. We defer the details of feature distillation for VLMs to CLIP-KD [101].

### v2

# Simple yet Effective Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

@@ Abstract
Semi-supervised learning (SSL) has emerged as a practical solution for addressing data scarcity challenges by leveraging unlabeled data. Recently, vision-language models (VLMs), pre-trained on massive image-text pairs, have demonstrated remarkable zero-/few-shot performance that often surpasses SSL approaches due to their exceptional generalization capabilities. This gap motivates us to question: how can we effectively harness the powerful generalization capabilities of VLMs into task-specific models? Knowledge distillation (KD) offers a natural framework for transferring VLM capabilities, but we identify that it suffers from gradient conflicts between supervised and distillation losses. To address this challenge, we propose Dual-Head Optimization (DHO), which introduces dual prediction heads for each distinct signal. We observe that DHO resolves gradient conflicts, enabling improved feature learning compared to single-head KD baselines, with practical benefits of minimal computational overhead and test-time hyperparameter tuning without retraining. Extensive experiments across 15 datasets show that DHO consistently outperforms KD baselines, often outperforming teacher models with smaller student models. DHO also achieves new state-of-the-art performance on both in-distribution ImageNet semi-supervised learning and out-of-distribution generalization across ImageNet variants. We publicly release our code and model checkpoints to facilitate future research at https://github.com/erjui/DHO.
@@ 1 Introduction
Vision-language models (VLMs), which learn joint vision-language representations through large-scale pre-training, have shown remarkable zero-shot capabilities across diverse tasks (Radford et al., 2021; Jia et al., 2021). Building upon these strong foundational capabilities, recent work has explored various adaptation strategies, including parameter-efficient approaches such as linear probing (Li et al., 2022; Huang et al., 2024), lightweight adapters (Zhang et al., 2021; Gao et al., 2024), and prompt-based fine-tuning methods (Zhou et al., 2022b; a; Lafon et al., 2025), demonstrating the potential of VLMs for data-limited visual recognition tasks.
In parallel, semi-supervised learning (SSL) has emerged as a practical approach to address data scarcity by leveraging both labeled and unlabeled data (Sohn et al., 2020; Assran et al., 2021; Cai et al., 2022; Zheng et al., 2023). While these methods have shown success to leverage large amounts of unlabeled data, they often struggle to match the impressive zero- and few-shot capabilities of large pre-trained VLMs (Liu et al., 2023b). This discrepancy highlights a fundamental limitation: traditional semi-supervised methods, despite their theoretical appeal, remain suboptimal compared to the rich representations learned by foundation models through massive-scale pre-training.
VLMs excel at zero- and few-shot generalization but may lack the fine-grained discriminative power needed for specific tasks, while models trained on limited labeled data capture task-specific patterns but generalize poorly. This complementary nature motivates us to integrate generalist VLM knowledge with task-specific supervision in semi-supervised learning settings. Therefore, the challenge naturally arises: how can we effectively transfer the powerful capabilities of large VLMs to task-specific models in semi-supervised settings?
@@
Knowledge distillation (KD; Hinton, 2015) emerges as a natural solution to the challenge, as it offeres an efficient framework to transfer knowledge from large VLMs to student models while simultaneously leveraging task-specific patterns from limited labeled data. However, existing VLM distillation methods primarily focus on general-purpose training (Yang et al., 2024a; Vasu et al., 2024; Udandarao et al., 2024; Yang et al., 2024c) or employ multi-stage pipelines with unsupervised pre-distillation (Vemulapalli et al., 2024; Wu et al., 2025), which require additional task-specific fine-tuning. Conventional single-stage KD methods, such as logit distillation (Hinton, 2015; Chen et al., 2020), offer a more direct approach, but we find them suboptimal in semi-supervised settings.
Through analysis, we identify that this stems from a fundamental problem: gradient conflicts between the supervised loss (from limited labeled data) and the distillation loss (from teacher predictions). This misalignment is particularly severe in semi-supervised settings, where the strong and consistent distillation signal from the teacher can overwhelm the weak and potentially noisy signal from scarce labeled data. Such gradient conflicts ( and in Fig.˜3) are well-documented to impair effective feature learning (Yu et al., 2020a; Liu et al., 2021; Chen & Er, 2025), preventing the model from finding an optimal balance between task-specific adaptation and general knowledge transfer.
@@
To address the above issue, we propose a simple yet effective distillation framework, DHO (Dual-Head Optimization), which jointly leverages labeled samples and the probabilistic outputs of the teacher model. Specifically, it learns two distinct heads, each optimized with a separate loss: the supervised and the KD loss, respectively. Our analysis reveals that DHO mitigates gradient conflicts both in the classification head and the shared feature extractor ( in Fig.˜3), arising from the two different training signals and improves feature representations compared to baselines (Sections˜3.3 and 8). Our framework also enables controlling the relative influence of supervised and teacher predictions through linear combination of both head outputs, whose effectiveness is demonstrated in Fig.˜6. Additionally, DHO enables post-training hyperparameter adjustment by tuning the linear combination weights at inference (Fig.˜9), eliminating costly training-time hyperparameter search.
We conduct extensive experiments across 15 different datasets including ImageNet (Russakovsky et al., 2015). The experimental results demonstrate the consistent improvement of DHO over conventional KD methods across all evaluated datasets (Fig.˜1-Left). DHO sometimes even outperforms the zero/few-shot teacher models with smaller student models with the same labeled data, demonstrating effective task-specific enhancement beyond teacher capabilities (Figs.˜4 and 5). Furthermore, DHO achieves new state-of-the-art (SoTA) performance on ImageNet semi-supervised setting, improving accuracy by 3% and 0.1% with 1% and 10% labeled data, respectively, while using fewer parameters (Fig.˜1-Right). Notably, DHO can be seamlessly integrated with existing adaptation techniques with minimal computational overhead, achieving new SoTA on Out-of-Distribution (OOD) tasks across ImageNet distribution-shifted variants (Tab.˜2 and Section˜E.5).
@@
  • We firstly identify gradient conflict when integrating VLMs’ general knowledge with task-specific supervision from limited data. To address this, we propose Dual-Head Optimization (DHO), which optimizes the supervised and distillation objectives in separate heads.
  • DHO effectively resolves gradient conflicts both in the classification head and the shared feature extractor, leading to improved feature representations. Our framework enables flexible post-training adjustment of dual head output weights with minimal computational overhead.
  • In extensive experiments on 15 datasets, DHO consistently outperforms baselines and sometimes surpasses teacher model performance. It establishes new SoTA for in-distribution ImageNet (e.g., +3%/+0.1% at 1%/10% labels) with fewer parameters (76M/767M), and also achieves SoTA OOD generalization across ImageNet variants when integrated with adaptation methods.
2 Method
2.1 Preliminaries
We begin with preliminaries: a brief background on VLMs, the problem formulation for few-/low-shot learning, and single-head KD baselines. We defer related work—vision–language pretraining, data-limited adaptation of VLMs, knowledge distillation, and dual-head methods—to Appendix˜A.
@@ 2.1 Preliminaries
[…]
@@
We focus on transferring knowledge from VLMs to task-specific models under few-shot or low-shot semi-supervised learning scenarios, where both labeled and unlabeled data are utilized. Specifically, given a $K$-shot and $C$-class classification problem, we are provided with a labeled dataset $\mathcal{D}^{(l)}=\{(x^{(l)}_{n},y_{n})\}_{n=1}^{N}$, where $N=K\times C$ is the total number of labeled examples, and $y_{n}\in\{1,\ldots,C\}$ denotes the class labels. Additionally, we have access to an unlabeled dataset $\mathcal{D}^{(u)}=\{x^{(u)}_{m}\}_{m=1}^{M}$ consisting of $M$ unlabeled images. Low-shot learning represents a more realistic setting than traditional few-shot learning where only a small fraction, e.g., 1% ($\frac{N}{N+M}\approx 0.01$) or 10% ($\frac{N}{N+M}\approx 0.1$) of the total dataset is labeled. Our goal is then to develop a student model by leveraging $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$, guided by the knowledge of the VLM encoders $f_{\mathcal{X}}$ and $f_{\mathcal{T}}$. The student model consists of a feature extractor $g(x)$ parameterzied by $\theta$ and a linear prediction head $h(z)=Wz+b\in\mathbb{R}^{c}$, where $W\in\mathbb{R}^{C\times d},b\in\mathbb{R}^{C}$, followed by the softmax function $\sigma$.
Figure 2: Conceptual illustration on KD frameworks, Single-Head Optimization (SHO) and Dual-Head Optimization (DHO), for semi-supervised settings. As demonstrated in Fig.˜3, we observe gradient conflict of SHO. In contrast, DHO mitigates such conflicts by leveraging dual-head architectures in Fig.˜3.
Single-head optimization (SHO) of KD.
Our method builds on logit distillation in semi-supervised settings (Hinton, 2015; Chen et al., 2020) that combines supervised loss $\mathcal{L}_{\text{CE}}$ on the labeled dataset $\mathcal{D}^{(l)}$ with KD loss $\mathcal{L}_{\text{KD}}$ on both labeled and unlabeled datasets $\mathcal{D}^{(l)}\cup\mathcal{D}^{(u)}$, i.e., $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$. Specifically, the supervised loss $\mathcal{L}_{\text{CE}}$ and the KD loss $\mathcal{L}_{\text{KD}}$ are defined as follows:
@@
$\mathcal{L}_{\text{KD}}$ $=\frac{1}{N}\sum_{n}D_{\mathrm{KL}}\left[p_{n}^{(l)}\|\sigma(h(z_{n}^{(l)}))\right]+\frac{1}{M}\sum_{m}D_{\mathrm{KL}}\left[p_{m}^{(u)}\|\sigma(h(z_{m}^{(u)}))\right].$ (2)
where $\ell$ denotes $D_{\mathrm{KL}}$ represent the cross-entropy and Kullback-Leibler divergence, respectively. $z_{n}^{(l)}=g(x_{n}^{(l)})$ and $z_{m}^{(u)}=g(x_{m}^{(u)})\in\mathbb{R}^{d}$ are feature representations obtained by the feature extractor $g$. $p_{n}^{(l)}$ and $p_{m}^{(u)}$ are the categorical probability vectors of labeled $x_{n}^{(l)}$ and unlabeled data $x_{m}^{(u)}$, respectively, obtained by teacher VLM encoders $f_{\mathcal{X}}$ and $f_{\mathcal{T}}$. Another well-studied single-head KD baseline is feature distillation, which leverages mean squared error (MSE) loss to directly align feature representations extracted by the student encoder $g$ and the teacher image encoder $f_{\mathcal{X}}$. We defer the details of feature distillation for VLMs to CLIP-KD (Yang et al., 2024a).